r=3 h=6 which of the following is equal to 2πr^2+2πrh

30π
45π
48π
54π
90π
1
Expert's answer
2014-01-16T10:04:57-0500
We just need to substitute the values r and h into equation.
2πr^2+2πrh=2π3^2+2π3*6=2π*9+2π*18=18π+36π=54π.
Right answer is 54π.